# thanks: equal contributionthanks: equal contribution

### v1

@@ VI Roadmap & Key Open Questions
Filamentous structures are ubiquitous in biology from cytoskeletal filaments to propulsive appendages to whole organisms. Despite spanning several orders of magnitude in scale and complexity, surprising parallels can be observed in how biological form translates into physiological function in these different contexts. In this review, we have sought to outline the core physical principles that govern these relationships and common frameworks for studying filamentous systems. As cutting-edge experimental technologies continue to deliver ever finer resolution of the molecular composition and architecture of the constituent parts of organelles, cells, and organisms, there is increasing need to integrate these insights with functional assays and biophysical models. At the same time, we should also be mindful that in many cases the natural context or habitat in which organisms live could also contribute to any observed phenomena Franks et al. (2016). Therefore, there is also a growing need to study and understand the behaviour of living systems ‘in the wild’, not just under controlled laboratory settings.
@@
Critical to many modelling and simulation approaches is detailed knowledge of the emergent material properties of the many types of filaments. Yet such measurements are not always available as they often rely on access to specialised technology, such as high-resolution imaging, atomic force microscopes, micropipette force sensors Backholm et al. (2013), or can only be performed on certain filaments, or in non-living systems or systems that have been rendered passive. This apparent disconnect requires more holistic, physics-inspired understanding of the context-dependent material properties of a filament and filamentous network, particularly due to their out-of-equilibrium activity. As a concrete example, the measurements of physical parameters such as Young’s modulus $E$ or the bending stiffness $\beta_{\mathrm{f}}$ of cytoskeletal filaments as shown in Table 1 strongly depends on the chemical and physical context.
@@
Recently, theories of active polymers Winkler and Gompper (2020) and active nematics Doostmohammadi et al. (2018) have identified interesting collective behaviour (such as bands, asters, and topological defects Liverpool et al. (2001); Duman et al. (2018); Bianco et al. (2018); Prathyusha et al. (2018); Vliegenthart et al. (2020)) that reproduce experiments with purified systems of microtubules and molecular motors Nédélec et al. (1997); Sanchez et al. (2012); Kruse et al. (2004); Sumino et al. (2012); Huber et al. (2018); Sciortino and Bausch (2021). While these systems shed light on the nonequilibrium dynamics of cytoskeletal filaments, the challenge that lies ahead is to develop these theories towards more realistic conditions like those inside the cell, which is a 3D crowded environment. Thus, these theoretical frameworks should be developed in tandem with experimental innovations. Identifying appropriate dimensionless quantities from filament properties, like the ones in Table 1, will enable the development of theories describing common phenomena in the collective dynamics of filaments. Doing so will allow the development of experimentally convenient proxy systems. Thus, efforts in characterising filaments, completing and expanding the data in Table 1 is required to develop a unified understanding of filament dynamics across scales.
@@
The study of active or self-propelled filaments in biology have led to advancement of new mathematical and physical theories, notably in the areas of topological braiding and entanglement, nonequilibrium statistical physics, and extreme mechanics such as supercoiling. However, one constraint in understanding filamentous systems is the practical challenge of acquiring good resolution data, and the lack of efficient computational tools to track filaments, particularly from 3D image stacks. There is also the open question of which metrics are most appropriate to describe the deformation of filaments and their collective organisation. Thus, devising universally-applicable, relatively system-independent measures for such systems requires further innovation.
@@
We speculate that, because living systems apparently increase entropy production Martyushev and Seleznev (2006); Davies et al. (2013) while producing more complex structures (Fig. 1) during their evolution, a generalised form of entropy might be a useful quantity in the search of system-independent measures. Because configurational degrees of freedom (positions, orientations) can be measured, it is possible to define structural complexity (or entropy) and use it to characterize living systems. Consider a function $\psi(\bm{x})$ representing the configuration of a system, for example, the atomic positions in a protein, the orientation of filaments in a network, or the muscle fibers in a vertebrate. The complexity of an ensemble can be quantified by means of a generalised Shannon–Boltzmann entropy Adami et al. (2000); Adami (2002); Bonchev and Buck (2005) ${H}=-\int D\psi\,P(\psi(\bm{x}))\ln P(\psi(\bm{x}))$, where $\int D\psi$ represents a path-integral sum over all possible configurations. We note that elongated shapes have five degrees of freedom (compare with the three degrees of freedom for a spherically symmetric object); furthermore, a flexible elongated shape has 5 degrees of freedom per unit length. Such explosion of degrees of freedom might constitute an entropic incentive for biological systems to evolve and use biological filaments.
It is also instructive to ask why filaments are a recurring motif across living systems. Not only do single units admit complex shape deformations, filaments can also easily self-assemble to form higher-order structures, or be reconfigured into networks of increasing complexity. These active physical structures have the capacity to perform complex functions without sophisticated control or centralised computation Wang et al. (2023); Moreau et al. (2024). Future applications, particularly in the realm of synthetic biology or bioengineering, may seek to harness the mechanical intelligence inherent in filamentous structures, to create novel adaptive materials.

### v2

@@ VI Roadmap & Key Open Questions
Filamentous structures are ubiquitous in biology from cytoskeletal filaments to propulsive appendages, and to whole organisms. Despite spanning several orders of magnitude in scale and complexity, surprising parallels can be observed in how biological form translates into physiological function. In this Review, we have sought to outline the core physical principles that govern these relationships and common frameworks for studying filamentous systems. As cutting-edge experimental technologies continue to deliver ever finer resolution of the molecular composition and architecture of the constituent parts of organelles, cells, and organisms, there is increasing need to integrate these insights with functional assays and biophysical models. At the same time, we should also be mindful that in many cases the natural context or habitat in which organisms live could also contribute to any observed phenomena [392]. Therefore, there is also a growing need to study and understand the behaviour of living systems ‘in the wild’, not just under controlled laboratory settings.
@@
Critical to many modelling and simulation approaches is detailed knowledge of the emergent material properties of the many types of filaments. Yet, such measurements are not always available as they often rely on access to specialised technology, such as high-resolution imaging, atomic force microscopes, micropipette force sensors [90], or can only be performed on certain filaments, or in non-living systems or systems that have been rendered passive. This apparent disconnect requires more holistic, physics-inspired understanding of the context-dependent material properties of a filament and filamentous network, particularly due to their out-of-equilibrium activity. As a concrete example, the measurements of physical parameters such as Young’s modulus $E$ or the bending stiffness $\beta_{\mathrm{f}}$ of cytoskeletal filaments as shown in Table 1 strongly depends on the chemical and physical context.
@@
Recently, theories of active polymers [68, 206] and active nematics [218] have identified interesting collective behaviour (such as bands, asters, and topological defects [393, 339, 394, 340, 395]) that reproduce experiments with purified systems of microtubules and molecular motors [216, 217, 396, 397, 398, 399]. While these systems shed light on the nonequilibrium dynamics of cytoskeletal filaments, the challenge that lies ahead is to develop these theories towards more realistic conditions like those inside the cell, which is a 3D crowded environment. Thus, these theoretical frameworks should be developed in tandem with experimental innovations. Identifying appropriate dimensionless quantities from filament properties (Table 1), will enable the development of theories describing common phenomena in the collective dynamics of filaments. Thus, efforts in characterising filaments, completing and expanding the data in Table 1 is required to develop a unified understanding of filament dynamics across scales.
Often, theoretical models developed to describe a specific system at one length scale can be applied across scales with little to no adaptation, as long as the mechanical properties of the filaments of interest are known. For instance, the active polymer model described in Eq. (11)–(13) has been successfully used to describe the dynamics of other filaments like flagella [213], cyanobacteria [214] or even worms [215]. The theory of nematics and their topological defects (see Eq. (14)-(15)) has been used to describe microtubules [223, 218], elongated bacteria [400, 401], and stem cells [402]. The mathematical description of metameric locomotion used for filamentous bacteria [309, 310] has also been suggested as potential description of the motion of annelids and myriapods [335]. Identifying commonalities and differences between filaments at different scales will allow the development of experimentally convenient proxy systems.
Beyond a consideration of possible shared tools, technologies and frameworks to analyse such complex spatiotemporal dynamics, the comparison begs the question of whether and to what extent the analytical or computational tools developed for modelling microscale active matter, can be adapted to study and predict dynamics at the macroscopic scale [372, 403]. Modified versions of the classic resistive force theory model for cylinders moving through fluids have been used to develop neuromechanical models of C. elegans crawling movement [404] with moderate empirical support [405], but the validity of these assumptions are yet to be fully demonstrated more generally for other organisms, particularly in the regime of $Re\sim 1$ where inertial effects are expected to play some role. Recent progress has been made on formulating slender body theories (originally developed for movement through Stokes flow) but applicable to larger anguilliform swimmers [406]. As we have already established, wave propagation is a particularly effective and universal means of breaking time-reversal symmetry and achieving self-propulsion across scales, yet the ultrastructure and biological mechanism of actuation can differ greatly across the organisms. Thus, it remains an open question how the same range of movement patterns can arise from completely different neuromuscular control architectures or even aneural mechanisms such as coordinated activity of molecular motors (in the case of cilia and flagella).
@@
The study of active or self-propelled filaments in biology have led to advancement of new mathematical and physical theories, notably in the areas of topological braiding and entanglement, nonequilibrium statistical physics, and extreme mechanics such as supercoiling. However, one constraint in understanding filamentous systems is the practical challenge of acquiring good resolution data, and the lack of efficient computational tools to track filaments, particularly from 3D image stacks. There is also the open question of which metrics are most appropriate to describe the deformation of filaments and their collective organisation; can we identify measures of effective activity that bridge various orders of magnitude in length scale? We think that devising universally-applicable, relatively system-independent measures for such systems requires further innovation.
@@
It is also instructive to return to the initial point of this Review and ask why filaments are a recurring motif across living systems? Mechanical forces are being increasingly recognised as shaping factors in biology [407], and geometry does create and influence mechanical interactions. We speculate that, because living systems increase entropy production [408, 409] while producing more complex structures (Fig. 1) during their evolution, a generalised form of entropy might be a useful quantity in the search of system-independent measures. Because configurational degrees of freedom (positions, orientations) can be measured, it is possible to define structural complexity (or entropy) and use it to characterize living systems. Consider a function $\psi(\bm{x})$ representing the configuration of a system, for example, the atomic positions in a protein, the orientation of filaments in a network, or the muscle fibers in a metazoan. The complexity of an ensemble can be quantified by means of a generalised Shannon–Boltzmann entropy [410, 411, 412] ${H}=-\int D\psi\,P(\psi(\bm{x}))\ln P(\psi(\bm{x}))$, where $\int D\psi$ represents a path-integral sum over all possible configurations. We note that slender rods have five degrees of freedom (compare with the three degrees of freedom for a spherically symmetric object); furthermore, a flexible elongated shape has five degrees of freedom per unit length. Such explosion of degrees of freedom might constitute an entropic incentive for biological systems to evolve and use biological filaments. Not only do single units admit complex shape deformations, filaments can also easily self-assemble to form higher-order structures, or be reconfigured into networks of increasing complexity. These active physical structures have the capacity to perform complex functions without sophisticated control or centralised computation [413, 414]. Future applications, particularly in the realm of synthetic biology or bioengineering, may seek to harness the mechanical intelligence inherent in filamentous structures, to create novel adaptive materials.
VII Acknowledgments
This work was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant 853560 EvoMotion, and UK Advanced Research and Invention Agency’s Nature Computes Better programme (K.Y.W). We also acknowledge support by UKRI/ Wellcome grant EP/T022000/1-PoLNET3 (J.C.). We thank an anonymous referee for helpful suggestions on Fig. 1 and Prof Alice Quillen for the use of the image in Fig. 5(d).